Using the method of substitution to help solve for this equation.
3x-2y=-4
2x-y=-4

Answer by ankor@dixie-net.com(22740)

Using the method of substitution to help solve for this equation.
3x - 2y = -4
2x - y = -4
add y and 4 to both sides and you have
2x + 4 = y
Replace y with (2x+4) in the first equation
3x - 2(2x+4) = -4
3x - 4x - 8 = -4
-x = -4 + 8
-x = 4
x has to be positive, multiply both sides by -1
x = -4
Find y
y = 2x + 4
y = 2(-4) + 4
y = -8 + 4
y = -4
:
Check in the first equation 3x - 2y = -4
3(-4) - 2(-4) = -4
-12 + 8 = -4